m + 3 divided by m - 3 < or equal to 2, D = {Reals}
Answer by ikdeep(226)

m + 3 divided by m - 3 < or equal to 2, D = {Reals}
It can be written as
(m + 3)/ m - 3 <= 2
multiply both sides by m-3 and we get
m + 3 <= 2m - 6
put together like variables and we get
0 <= 2m - m -6 - 3
0 <= m - 9
since D = real
therefore m-9 >= 0
subtract 9 from both sides,,we get

m >= 9
hope this will help you
Please feel free to revert back for any further queries.